OpenStudy (anonymous):

What is the equivalent equation to 5 - 3x^2 + 5x - 7 = 2x + 1? :D

OpenStudy (anonymous):

are you looking for the simplified form ?

OpenStudy (anonymous):

I don't know what it is they're asking for exactly. These are my option: 3x^2 - 3x + 3 = 0 -3x^2 - 3x + 13 = 0 3x^2 + 15x + 3 = 0 -3x^2 - 3x + 3 = 0

OpenStudy (anonymous):

yes, you need to simplify the equation..

OpenStudy (anonymous):

Mmkay so it's just basic simplifying or is their some extra ish that has to be done? lol

OpenStudy (anonymous):

I think it's b or d....

OpenStudy (anonymous):

lol, i got this...